OpenStudy (anonymous):

The defining equations for the function f is f(x) = 2x^2 + 3. The defining equation for the function g is g(x) = x-5. Then the defining equation for g \circle f is:

OpenStudy (anonymous):

18x^2 - 6x + 3 18x^2 + 6x + 3 2x^2-2 6x^3 - 10x^2 + 9x - 15

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@surd @wolfe8 @emmigrace222 @vivalatee @niaispretty @marylou004 @LazyBoy

OpenStudy (***[isuru]***):

\[(gof)(x) = g[f(x)]\] can u find g [ f(x)] ?

OpenStudy (anonymous):

how owuld i do that?

OpenStudy (***[isuru]***):

ok... let say f(x) = 2x^2 + 5x then do u know how to find f(3) ?

OpenStudy (anonymous):

nope im sorry i dont understand how to do thease

OpenStudy (***[isuru]***):

it's fine... if f(x) = 2x^2 + 5x and if they asks u to find f(3) all u have to do is substitute 3 to the places where x is in f(x).. in this case f(3) = 2(3)^2 + 5(3) got it ?

OpenStudy (anonymous):

so would i than figure out that proplem?

OpenStudy (***[isuru]***):

if u need to figure out the problem ... u need to understand the above thing.. basically if they give u an expression f(x) and asks u to find f(g) u have to substitute g instead of x in f(x)... in ur case u've got g(x) = x - 5 and u need to find g[f(x)] then u have to substitute f(x) instead of x in g(x)... g(x) = x -5 = f(x) - 5 and u know that f(x) = 2x^2 + 3 then u can write that value in place of f(x) in g[f(x)] g[f(x)] = f(x) - 5 = 2x^2 + 3 -5 = 2x^2 - 2 that's it and hope this will help ya out!!